Please show step by step, thank you!Calculate how many milliliters of 0.622 M KOH should be added to 4.90 g of MOPS (see the Structures and pKa Values for Common
Answer #1

Firstly, listing some vital values for MOPS from buffer's table:

IUPAC name: 3-morpholinopropane-1-sulphonic acid

Formula: C7H15NO4S

Molar mass: 209.2633 g/mol

pKa = 7.20

From the structure, it's clear that it is a mono-basic acid

----------------------------------------------------------------------------------------------------

Simply, this mono-basic acid can be represented by HA for the sake of this question.

Now, every one mole of KOH added will react with one mole of H+ ions derived from HA and reduce [ HA] or [acid] by 1 mole and increase the concentration of its conjugate base (A-) by 1 mole. See the following chemical equation:

HA(aq) + KOH(aq)\rightleftharpoons A-(aq) + H2O(l) + K+(aq)

Now calculating the number of moles of MOPS from 4.90 g

Number of moles of MOPS = (4.90 g)\times(1mol)/(209.2633g) = 0.0234 mol = 23.4 milli-moles

< Conversion of moles to milli-moles for convenience: 0.0234 mol = 0.0234 mol \times (1000 m moles)/1mol) = 23.4 milli moles>

Let the number of milli-moles of KOH added be 'x'.

Then, applying Henderson-Hasselbach's equation:

pH = pKa + log ([base+x]/[acid-x])

Filling in the values given as well as the number of milli-moles calculated above:

7.77 = 7.20 + log{(23.4+x)/23.4-x)}

=> log{(23.4+x)/23.4-x)} = 0.57

=> (23.4+x)/(23.4-x) = 100.57

=> (23.4+x)/(23.4-x) = 3.715

=> 23.4 + x = 86.931 - 3.715x

=> 4.715x = 86.931 - 23.4

=> 4.715x = 63.531

=> x = 13.474

\therefore 13.474 milli-moles of KOH must be added to get the pH of 7.77

Now, finding the volume in mL of KOH solution of 0.622mol/L that contains 13.474 milli-moles

Volume of KOH solution to be added = (13.474 \times 10-3 moles) \times (1000mL/0.622 moles) = 21.66 mL (Answer)